# Monotonicity of the ultrafilter number function

### v3

@@ II $q_{0}$ $q_{1}$cdots $q_{\omega}$ $q_{\omega+1}$cdots
$\mathbb{P}$ is $\alpha$-strategically closed if Player II has a winning strategy in the game $\Gamma_{\alpha}(\mathbb{P})$. If $\kappa$ is a cardinal and $\mathbb{P}$ is $\kappa$-strategically closed, then it is easy to see that $\mathbb{P}$ is $\kappa$-Baire, and the forcing with $\mathbb{P}$ does not add new $<\kappa$-sequences.
Fact 6.5.
Let $\kappa$ be a cardinal, and $\langle{\mathbb{P}_{n},\dot{\mathbb{Q}}_{n}\mid n<\omega}\rangle$ be an $\omega$-stage iteration such that each ˙$\dot{\mathbb{Q}}_{n}$ is forced to be $\kappa$-strategically closed. Then the inverse limit of the $\mathbb{P}_{n}$’s is $\kappa$-strategically closed.
For a regular uncountable cardinal $\mu$ and regular $\lambda<\mu$, let $\mathbb{S}_{\mu,\lambda}$ be the poset of all bounded subsets $p\subseteq\{\alpha<\mu\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ such that $p$ has a maximal element $\max(p)$ and for every $\beta\leq\max(p)$, $p\cap\beta$ is non-stationary in $\beta$. Define $p\leq q$ if $p$ is an end-extension of $q$.
@@
$\left|{\mathbb{S}}\right|=2^{<\mu}$, so $\mathbb{S}$ satisfies the $(2^{<\mu})^{+}$-c.c.
@@
    $\mathbb{S}$ is $\mu$-strategically closed, hence the forcing with $\mathbb{S}$ does not add new $<\mu$-sequences.
@@
(1). For a limit $\beta<\mu$ and a partial play $\langle p_{\xi},q_{\zeta}\mid\xi,\zeta<\beta$, $\xi=0$ or $\rangle$, suppose $q=\bigcup_{\zeta<\beta}q_{\zeta}$ is non-reflecting. Let $\gamma=\sup(q)$. Then Player II takes $q\cup\{\gamma+\lambda\}$ as his move. This is a winning strategy of Player II.
(2). Non-reflectingness is immediate from the definition of $\mathbb{S}$. To show that $\bigcup G$ is stationary, take a name $\dot{C}$ for a club in $\mu$. Take a descending sequence $\langle{p_{i}\mid i<\lambda}\rangle$ such that for every $i<\lambda$, there is $\alpha_{i}$ such that $\alpha_{i}<\sup(p_{i})<\alpha_{i+1}$ and $p_{i+1}\Vdash\check{\alpha}_{i}\in\dot{C}$. Let $\alpha=\sup_{i}\alpha_{i}$, and $p=\bigcup_{i<\lambda}p_{i}\cup\{\alpha\}$. It is easy to check that $p\in\mathbb{S}$ and $p\Vdash\alpha\in\dot{C}\cap\bigcup\dot{G}$. ∎
@@
Let $\kappa$ be a singular cardinal with countable cofinality, and suppose GCH. Let $\langle{\kappa_{n}\mid n<\omega}\rangle$ be a strictly increasing sequence of regular uncountable cardinals with limit $\kappa$, and $\lambda$ a regular uncountable cardinal with $\lambda<\kappa_{0}$. Then there is a poset $\mathbb{P}$ such that $\mathbb{P}$ preserves all cardinals, and adds a sequence $\langle{S_{n}\subseteq\kappa_{n}\mid n<\omega}\rangle$ satisfying the conditions of Proposition 6.2.
@@
Fix such $\langle{\kappa_{n}\mid n<\omega}\rangle$ and $\lambda$. Define $\mathbb{P}_{n}$ and ˙$\dot{\mathbb{Q}}_{n}$ for $n<\omega$ as follows. Let $\mathbb{P}_{0}$ be the trivial poset. Suppose $\mathbb{P}_{n}$ is defined. Then ˙$\dot{\mathbb{Q}}_{n}$ is a $\mathbb{P}_{n}$-name such that
@@
Put $\mathbb{P}_{n+1}$ as the forcing product $\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}$. It is routine to show that $\left|{\mathbb{P}_{n+1}}\right|\leq 2^{<\kappa_{n}}=\kappa_{n}$, so it satisfies the $(\kappa_{n})^{+}$-c.c. Let $\mathbb{P}$ be the inverse limit of the $\mathbb{P}_{n}$’s. We check that $\mathbb{P}$ is a required poset.
By Fact 6.5, for each $n<\omega$, $\mathbb{P}$ can be factored as a forcing product $\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}*\dot{\mathbb{P}}_{tail}$ such that $\Vdash_{\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}}$˙$\dot{\mathbb{P}}_{tail}$ is $\kappa_{n+1}$-strategically closed”. This factorization shows that the forcing with $\mathbb{P}$ preserves all cofinalities.
Take a $(V,\mathbb{P})$-generic $G$, and let $S_{n}$ be a generic subset of $\kappa_{n}$ induced by $G$. By the factorization before, we have that $S_{n}$ is a non-reflecting stationary subset of $\{\alpha<\kappa_{n}\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ in $V[G]$.
Finally, we see that the sequence $\langle{S_{n}\mid n<\omega}\rangle$ satisfies the the conditions of Proposition 6.2. To do this, return to $V$. Let $\theta$ be a sufficiently large regular cardinal and $N=H_{\theta}$. Let $\dot{S}_{n}$ be a canonical name for $S_{n}$. Take a $\mathbb{P}$-name $\dot{f}$ such that $p\Vdash\dot{f}:[\check{N}]^{<\omega}\to N$. It is enough to show that there is $X\subseteq N$ and $q\leq p$ such that $\left|{X}\right|=\lambda\subseteq X$, $[X]^{\omega}\subseteq X$, and
$q\Vdash\dot{f}``[\check{X}]^{<\omega}\subseteq\check{X},\sup(\check{X}\cap\kappa_{n})\in\dot{S}_{n}\text{ for $n<\omega$}.$
Take a large regular cardinal $\chi>\theta$ and $M\prec H_{\chi}$ containing all relevant objects such that $\left|{M}\right|=\lambda\subseteq M$ and $[M]^{<\lambda}\subseteq M$. Let $X=M\cap H_{\theta}$. Note that $\left|{X}\right|=\lambda$, $[X]^{\omega}\subseteq X$, and $\mathord{\mathrm{cf}}(\sup(X\cap\kappa_{n}))=\lambda$ for every $n<\omega$. Fix an enumeration $\langle{x_{i}\mid i<\lambda}\rangle$ of all elements of $X$. By the $\lambda$-strategically closedness of $\mathbb{P}$ and the closure property of $M$, we can take a descending sequence $\langle{q_{i}\mid i<\lambda}\rangle$ in $\mathbb{P}$ such that the following holds:
@@
    For $i<\lambda$, there is $y_{i}\in X$ such that $q_{i}\Vdash\dot{f}(\check{x}_{i})=\check{y}_{i}$.
@@
    For $i<\lambda$ and $n<\omega$, there is $\alpha\in X\cap\kappa_{n}$ such that $q_{i+1}\Vdash\max(q_{i}(n))=\check{\alpha}$.
Then take a function $q$ on $\omega$ as follows: $q(n)$ is a $\mathbb{P}_{n}$-name such that $\Vdash_{\mathbb{P}_{n}}$$q(n)=\bigcup_{i<\lambda}q_{i}(n)\cup\{\sup(\check{X}\cap\check{\kappa}_{n})\}$. It is straightforward to show that $q$ is a lower bound of the $q_{i}$’s, and $q\Vdash\dot{f}``[\check{X}]^{<\omega}\subseteq\check{X},\sup(\check{X}\cap\kappa_{n})\in\dot{S}_{n}$ for $n<\omega$. ∎
By this proposition, if $\kappa$ is measurable we can take a generic extension in which $\kappa$ is measurable and there is no $\omega_{1}$-indecomposable uniform ultrafilter over $\omega_{\omega}$.
@@
(1). Take a $(V,\mathrm{Add}(\omega,\mu))$-generic $G$. In $V[G]$, we have $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{1})$. We claim that, in $V[G]$, $\omega_{\omega}$ does not have an $\omega_{1}$-indecomposable uniform ultrafilter. Suppose to the contrary that there is an $\omega_{1}$-indecomposable uniform ultrafilter $\mathcal{U}$. Let $\dot{\mathcal{U}}$ be a name for $\mathcal{U}$. In $V$, let =$\mathcal{F}=\{X\subseteq\omega_{\omega}\mid\Vdash X\in\dot{\mathcal{U}}\}$. $\mathcal{F}$ is a uniform filter over $\omega_{\omega}$, and since $\mathrm{Add}(\omega,\mu)$ has the c.c.c. and $\mathcal{U}$ is $\omega_{1}$-indecomposable, for every $f:\omega_{\omega}\to\omega_{1}$ there is $\gamma<\omega_{1}$ such that $\{\alpha\mid f(\alpha)\leq\gamma\}\in\mathcal{F}$. Then every ultrafilter extending $\mathcal{F}$ is an $\omega_{1}$-indecomposable uniform ultrafilter, this is a contradiction.
Now, in $V[G]$, every uniform ultrafilter over $\omega_{\omega}$ is $\omega_{1}$-decomposable. By Lemma 5.2, we have that $\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\omega_{\omega})$, hence $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\omega_{\omega})$.
@@
We will show that both $\mathfrak{u}(\kappa^{+\omega})<\mathfrak{u}(\omega_{\omega})$ and $\mathfrak{u}(\kappa^{+\omega_{1}})<\mathfrak{u}(\omega_{\omega})$ are consistent as well. See Corollaries 7.7 and Proposition 7.18.

### v4

@@ II $q_{0}$ $q_{1}$ $\cdots$ $q_{\omega}$ $q_{\omega+1}$ $\cdots$
$\mathbb{P}$ is $\alpha$-strategically closed if Player II has a winning strategy in the game $\Gamma_{\alpha}(\mathbb{P})$. If $\kappa$ is a cardinal and $\mathbb{P}$ is $\kappa$-strategically closed, then forcing with $\mathbb{P}$ does not add new $\kappa-$sequences.
Proposition 6.5.
Let $\kappa$ be a cardinal, and $\langle{\mathbb{P}_{n},\dot{\mathbb{Q}}_{n}\mid n<\omega}\rangle$ be an $\omega$-stage iteration such that $\Vdash_{\mathbb{P}_{n}}$“ $\dot{\mathbb{Q}}_{n}$ is $\kappa$-strategically closed” for every $n<\omega$. Then the inverse limit of the $\mathbb{P}_{n}$’s is $\kappa$-strategically closed.
For a regular uncountable cardinal $\mu$ and regular $\lambda<\mu$, let $\mathbb{S}_{\mu,\lambda}$ be the poset of all bounded subsets $p\subseteq\{\alpha<\mu\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ such that $p$ has a maximal element $\max(p)$ and for every $\beta\leq\max(p)$ with $\mathord{\mathrm{cf}}(\beta)>\omega$, $p\cap\beta$ is non-stationary in $\beta$. Define $p\leq q$ if $p$ is an end-extension of $q$.
@@
$\left|{\mathbb{S}}\right|=2^{<\mu}$, so $\mathbb{S}$ satisfies the $(2^{<\mu})^{+}$-$c.$
@@
    $\mathbb{S}$ is $\lambda^{+}$-closed.
  2. (2)
    $\mathbb{S}$ is $\mu$-strategically closed, hence forcing with $\mathbb{S}$ does not add new $\mathop{<}\mu$-sequences.
@@
(1). For the $\lambda^{+}$-closedness, take a descending sequence $\langle{p_{\xi}\mid\xi<\alpha}\rangle$ for some $\alpha<\lambda^{+}$. Let $\gamma=\sup_{\xi<\alpha}(\max(p_{\xi}))$. We know $\mathord{\mathrm{cf}}(\gamma)\leq\lambda$, hence $\bigcup_{\xi<\alpha}p_{\xi}$ is non-stationary in $\gamma$. Let $p=\bigcup_{\xi<\alpha}p_{\xi}\cup\{\gamma+\lambda\}$. we can check that $p\in\mathbb{S}$ and $p$ is a lower bound of the $p_{\xi}$’s.
[…]
(3). Non-reflectingness is immediate from the definition of $\mathbb{S}$. To show that $\bigcup G$ is stationary, take a name $\dot{C}$ for a club in $\mu$. Take a descending sequence $\langle{p_{i}\mid i<\lambda}\rangle$ such that for every $i<\lambda$, there is $\alpha_{i}$ such that $\alpha_{i}<\max(p_{i})<\alpha_{i+1}$ and $p_{i+1}\Vdash\check{\alpha}_{i}\in\dot{C}$. Let $\alpha=\sup_{i}\alpha_{i}$, and $p=\bigcup_{i<\lambda}p_{i}\cup\{\alpha\}$. It is easy to check that $p\in\mathbb{S}$ and $p\Vdash\alpha\in\dot{C}\cap\bigcup\dot{G}$. ∎
@@
Let $\kappa$ be a singular cardinal with countable cofinality, and suppose GCH. Let $\langle{\kappa_{n}\mid n<\omega}\rangle$ be a strictly increasing sequence of regular uncountable cardinals with limit $\kappa$, and $\lambda$ a regular uncountable cardinal with $\lambda<\kappa_{0}$. Then there is a poset $\mathbb{P}$ such that $\left|{\mathbb{P}}\right|=\kappa^{+}$, $\mathbb{P}$ preserves all cardinals, and adds a sequence $\langle{S_{n}\subseteq\kappa_{n}\mid n<\omega}\rangle$ satisfying the conditions of Proposition 6.2.
@@
Fix such $\langle{\kappa_{n}\mid n<\omega}\rangle$ and $\lambda$. Define $\mathbb{P}_{n}$ and $\dot{\mathbb{Q}}_{n}$ for $n<\omega$ as follows. Let $\mathbb{P}_{0}$ be the trivial poset. Suppose $\mathbb{P}_{n}$ is defined. Then $\dot{\mathbb{Q}}_{n}$ is a $\mathbb{P}_{n}$-name such that
@@
Let $\mathbb{P}_{n+1}=\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}$. It is routine to show that $\left|{\mathbb{P}_{n+1}}\right|\leq 2^{<\kappa_{n}}=\kappa_{n}$, so it satisfies the $(\kappa_{n})^{+}$-c.c. Let $\mathbb{P}$ be the inverse limit of the $\mathbb{P}_{n}$’s. We check that $\mathbb{P}$ witnessing the theorem.
A standard argument shows $\left|{\mathbb{P}}\right|=\kappa^{\omega}=\kappa^{+}$ and $\mathbb{P}$ is $\lambda^{+}$-closed. By Proposition 6.5, for each $n<\omega$, $\mathbb{P}$ can be factored as a forcing product $\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}*\dot{\mathbb{P}}_{tail}$ such that $\Vdash_{\mathbb{P}_{n}*\dot{\mathbb{Q}}_{n}}$$\dot{\mathbb{P}}_{tail}$ is $\kappa_{n+1}$-strategically closed”. This factorization shows that forcing with $\mathbb{P}$ preserves all cofinalities $\mathop{<}\kappa$. Since $\left|{\mathbb{P}}\right|=\kappa^{+}$, it preserves all cofinalities $\kappa^{+}$. We show that $\mathbb{P}$ preserves $\kappa^{+}$. If $(\kappa^{+})^{V}$ is collapsed in $V^{\mathbb{P}}$, since $\kappa$ is singular, the cofinality of $(\kappa^{+})^{V}$ is $\mathop{<}\kappa$ in $V^{\mathbb{P}}$. Hence there is a cofinal map $f$ from some $\alpha<\kappa$ into $(\kappa^{+})^{V}$. By the above factorization, this $f$ was added by $\mathbb{P}_{n}$ for some $n<\omega$, so $(\kappa^{+})^{V}$ is collapsed in $V^{\mathbb{P}_{n}}$. This is impossible because the cardinality of $\mathbb{P}_{n}$ is $\kappa$.
Take a $(V,\mathbb{P})$-generic $G$, and let $S_{n}$ be a generic subset of $\kappa_{n}$ induced by $G$. By the above factorization, we have that $S_{n}$ is a non-reflecting stationary subset of $\{\alpha<\kappa_{n}\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ in $V[G]$.
Finally, we see that the sequence $\langle{S_{n}\mid n<\omega}\rangle$ satisfies the the conditions of Proposition 6.2. To do this, return to $V$. Let $\theta$ be a sufficiently large regular cardinal and $N=H_{\theta}$. Let $\dot{S}_{n}$ be a canonical name for $S_{n}$. Take a $\mathbb{P}$-name $\dot{f}$ such that $p\Vdash\dot{f}:[\check{N}]^{<\omega}\to\check{N}$. It is enough to show that there is $X\subseteq N$ and $q\leq p$ such that $\left|{X}\right|=\lambda\subseteq X$, $^{\omega}X\subseteq X$, and
$q\Vdash\dot{f}``[\check{X}]^{<\omega}\subseteq\check{X},\sup(\check{X}\cap\check{\kappa}_{n})\in\dot{S}_{n}\text{ for $n<\omega$}.$
Take a large regular cardinal $\chi>\theta$ and $M\prec H_{\chi}$ containing $N,p,\langle{\dot{S}_{n}\mid n<\omega}\rangle,\dot{f}$ and other relevant objects such that $\left|{M}\right|=\lambda\subseteq M$ and $^{<\lambda}M\subseteq M$. Let $X=M\cap H_{\theta}=M\cap N$. Note that $\left|{X}\right|=\lambda$, $^{\omega}X\subseteq X$, and $\mathord{\mathrm{cf}}(\sup(X\cap\kappa_{n}))=\lambda$ for every $n<\omega$.
Since $\mathbb{P}$ is $\lambda^{+}$-closed, $\left|{M}\right|=\lambda\subseteq M$, and $^{<\lambda}M\subseteq M$, it is easy to take a descending sequence $\langle{q_{i}\mid i<\lambda}\rangle$ in $\mathbb{P}$ such that:
@@
    For every dense set $D\in M$ in $\mathbb{P}$, there is $i<\lambda$ with $q_{i}\in D\cap M$.
By the condition (2) and the elementarity of $M$, we have:
@@
    For every $x\in[M\cap N]^{<\omega}$, there is $i<\lambda$ and $y\in M\cap N(=X)$ such that $q_{i}\Vdash\dot{f}(\check{x})=\check{y}$.
  2. (4)
    For $i<\lambda$ and $n<\omega$, there is $j<\lambda$ and $\alpha_{i,n}\in M\cap N\cap\kappa_{n}(=X\cap\kappa_{n})$ such that $q_{j}\restriction n\Vdash_{\mathbb{P}_{n}}\max(q_{i}(n))=\check{\alpha}_{i,n}$.
Claim 6.9.
$\sup(X\cap\kappa_{n})=\sup_{i<\lambda}\alpha_{i,n}$ for every $n<\omega$.
Proof.
Since $\alpha_{i,n}\in M\cap N\cap\kappa_{n}=X\cap\kappa_{n}$, we have $\sup(X\cap\kappa_{n})\geq\sup_{i<\lambda}\alpha_{i,n}$. For the converse, take $\alpha\in X\cap\kappa_{n}$. Then the set $D=\{r\in\mathbb{P}\mid r\restriction n\Vdash_{\mathbb{P}_{n}}\max(r(n))\geq\check{\alpha}\}$ is dense in $\mathbb{P}$ and is in $M$. By the condition (4) there is $i<j<\lambda$ such that $q_{i}\in M\cap D$ and $q_{j}\restriction n\Vdash_{\mathbb{P}_{n}}\max(q_{i}(n))=\check{\alpha}_{i,n}$. Thus $\alpha_{i,n}\geq\alpha$, so $\sup(X\cap\kappa_{n})\leq\sup_{i<\lambda}\alpha_{i,n}$. ∎
Take a function $q$ on $\omega$ as follows: $q(n)$ is a $\mathbb{P}_{n}$-name such that $\Vdash_{\mathbb{P}_{n}}$$q(n)=\bigcup_{i<\lambda}q_{i}(n)\cup\{\sup(\check{X}\cap\check{\kappa}_{n})\}$. Finally we show that $q$ is a lower bound of the $q_{i}$’s, and $q\Vdash\dot{f}``[\check{X}]^{<\omega}\subseteq\check{X},\sup(\check{X}\cap\check{\kappa}_{n})\in\dot{S}_{n}$ for $n<\omega$.
Claim 6.10.
$q\in\mathbb{P}$ and $q$ is a lower bound of $\{q_{i}\mid i<\lambda\}$.
Proof.
It is enough to show that $q\restriction n\in\mathbb{P}_{n}$ and $q\restriction n$ is a lower bound of $\{q_{i}\restriction n\mid i<\lambda\}$ for every $n<\omega$. We prove this by induction on $n<\omega$. The case $n=0$ is trivial. Suppose $q\restriction n\in\mathbb{P}_{n}$ and is a lower bound of $\{q_{i}\restriction n\mid i<\lambda\}$. Then $q\restriction n\Vdash_{\mathbb{P}_{n}}$$\{q_{i}(n)\mid i<\lambda\}$ is a descending sequence in $\dot{\mathbb{Q}}_{n}=\dot{\mathbb{S}}_{\kappa_{n},\lambda}$”. By the condition (4) and the previous claim, we know that $q\restriction n\Vdash_{\mathbb{P}_{n}}$$\sup(\{\max(q_{i}(n))\mid i<\lambda\})=\sup(\check{X}\cap\check{\kappa}_{n})$”. Since $\mathord{\mathrm{cf}}(X\cap\kappa_{n})=\lambda$, we also have $q\restriction n\Vdash_{\mathbb{P}_{n}}$$\bigcup_{i<\lambda}q_{i}(n)$ is non-stationary in $\sup(\check{X}\cap\check{\kappa}_{n})$”. These observations show that $q\restriction\Vdash_{\mathbb{P}_{n}}$$q(n)=\bigcup_{i<\lambda}q_{i}(n)\cup\{\sup(\check{X}\cap\check{\kappa}_{n})\}\in\dot{\mathbb{Q}}_{n}$”, so $q\restriction(n+1)\in\mathbb{P}_{n+1}$. By the definition of $q$, we also have $q\restriction(n+1)$ is a lower bound of $\{q_{i}\restriction(n+1)\mid i<\lambda\}$. ∎
Now, since $q$ is a lower bound of $\{q_{i}\mid i<\lambda\}$, we know that $q\Vdash\dot{f}``[\check{X}]^{<\omega}\subseteq\check{X}$ by the condition (3). Moreover $q\Vdash\sup(\check{X}\cap\check{\kappa}_{n})\in\dot{S}_{n}$ for every $n<\omega$. This completes the proof.
By this proposition, we can construct a model in which there is a measurable cardinal and there is no $\omega_{1}$-indecomposable uniform ultrafilter over $\omega_{\omega}$; Suppose GCH holds and there is a measurable cardinal $\nu$. Let $\mathbb{P}$ be a poset from Proposition 6.8 for $\kappa=\omega_{\omega}$ and $\lambda=\omega_{1}$. In $V^{\mathbb{P}}$ there is no $\omega_{1}$-indecomposable uniform ultrafilter over $\omega_{\omega}$ by Proposition 6.2. The cardinality of $\mathbb{P}$ is $\mathop{<}\nu$, hence $\nu$ remains a measurable cardinal in $V^{\mathbb{P}}$ by Levy-Solovay Theorem (e.g., see Proposition 10.15 in Kanamori [15]).
@@
(1). Take a $(V,\mathrm{Add}(\omega,\mu))$-generic $G$. In $V[G]$, we have $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{1})$ by Theorem 4.5. We claim that, in $V[G]$, $\omega_{\omega}$ does not have an $\omega_{1}$-indecomposable uniform ultrafilter. Suppose to the contrary that there is an $\omega_{1}$-indecomposable uniform ultrafilter $\mathcal{U}$. Let $\dot{\mathcal{U}}$ be a name for $\mathcal{U}$. In $V$, let $\mathcal{F}=\{X\subseteq\omega_{\omega}\mid\,\Vdash\check{X}\in\dot{\mathcal{U}}\}$. $\mathcal{F}$ is a uniform filter over $\omega_{\omega}$, and since $\mathrm{Add}(\omega,\mu)$ has the c.c.c. and $\mathcal{U}$ is $\omega_{1}$-indecomposable, for every $f:\omega_{\omega}\to\omega_{1}$ there is $\gamma<\omega_{1}$ such that $\{\alpha\mid f(\alpha)\leq\gamma\}\in\mathcal{F}$. This shows that $\mathcal{F}$ is $\omega_{1}$-indecomposable in $V$. By Lemma 2.3, every ultrafilter extending $\mathcal{F}$ is an $\omega_{1}$-indecomposable uniform ultrafilter, this is a contradiction.
Now, in $V[G]$, every uniform ultrafilter over $\omega_{\omega}$ is $\omega_{1}$-decomposable. By Corollary 5.2, we have that $\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\omega_{\omega})$, hence $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\omega_{\omega})$.
@@
We will show that both $\mathfrak{u}(\kappa^{+\omega})<\mathfrak{u}(\omega_{\omega})$ and $\mathfrak{u}(\kappa^{+\omega_{1}})<\mathfrak{u}(\omega_{\omega})$ are consistent as well. See Corollary 7.7 and Proposition 7.17.